# %****␣bookchapter.tex␣Line␣75␣****https://informatics.tuwien.ac.at/

### v2

@@ 1 Introduction
Brain development during the first few months of life is a period of rapid structural and functional reorganization, making it a critical window for identifying potential neurodevelopmental deficits. Accurately predicting developmental outcomes during this period is essential for enabling early interventions that can mitigate the lifelong impact of developmental delays. Neonatal fMRI data, such as those from the Developing Human Connectome Project (dHCP), have shown potential for predicting neurodevelopmental outcomes [1] [2] [3] [4]. However, the spatiotemporal complexity of neonatal brain activity presents significant challenges for conventional analysis methods. This study investigates the potential of the Swin 4D fMRI Transformer (SwiFT) [19], a deep learning architecture designed to process high-dimensional fMRI data, for predicting neurodevelopmental outcomes from neonatal fMRI. Unlike existing methods, SwiFT leverages 4D spatiotemporal attention mechanisms to effectively capture dynamic brain connectivity patterns, offering a novel approach for analyzing neonatal fMRI data. Specifically, we aim to predict Bayley Scales of Infant and Toddler Development, Third Edition (Bayley-III/BSID-III) composite scores, encompassing cognitive, language, and motor skills, using neonatal fMRI from the dHCP dataset. To address the challenges of limited neonatal data and high dimensionality, we explore dimensionality reduction using group Independent Component Analysis (ICA) and pretraining SwiFT on large publicly available adult fMRI datasets. We hypothesize that (1) SwiFT outperforms baseline models in predicting Bayley-III scores due to its ability to capture complex spatiotemporal patterns in fMRI data and (2) incorporating ICA features and pretraining further improves predictive accuracy. Furthermore, we aim to interpret SwiFT’s predictions using Integrated Gradients with Smoothgrad sQuare (IG-SQ) [28] [29] [30] to identify brain regions associated with early cognitive, language, and motor development. This research advances the state of neurodevelopmental prediction by introducing a robust and interpretable framework for neonatal fMRI analysis. By addressing key methodological challenges, our findings lay the groundwork for future applications in neurodevelopmental research and clinical practice.
In Section 2, we introduce the dataset, its acquisition, (additional) processing and target values. We also explain SwiFT, our motivations and how we will interpret the results of the model. In Section 3, we will explain in what settings we perform our experiments, and introduce (i) the baselines, (ii) SwiFT running on (pre-processed) raw fMRI data with single-label and multi-label prediction, and (iii) SwiFT running on group-ICA-extracted features with single-label and multi-label prediction. In the end, we give an overview of the performance gain through our methods. In Section 4, we perform IG-SQ on the best performing model and provide our interpretation of the IG maps on risk of delay in cognitive, language and motor skills. Moreover, we provide a small glimpse into what model interpretations on brain-network-level may look like by using ICA-extracted features. Lastly, we conclude our research in Section 5 through discussing future directions and limitations of our current work.
@@ 2 Methods
Following prior studies  [10] [11], we simplify the classification task into a binary format, with a threshold of 85 distinguishing low- and high-risk cases. Regression tasks are also conducted to predict continuous Bayley-III scores.
@@
To predict neurodevelopmental outcomes from neonatal fMRI, we employed SwiFT, a Transformer-based architecture designed to model high-dimensional spatiotemporal fMRI data. SwiFT processes raw 4D fMRI volumes, capturing local and global dependencies critical for neurodevelopmental predictions. The model operates by dividing 4D fMRI data into patches, embedding them as tokens, and processing these tokens through multiple layers of 4D Swin Transformer blocks. These blocks utilize a shifted window multi-head self-attention mechanism, enabling efficient modeling of complex spatiotemporal patterns while reducing computational complexity through patch merging. Previously, SwiFT demonstrated robust performance on diverse fMRI tasks, including sex classification (97.7% accuracy on HCP) and age prediction (MAE of 3.40 years on UK Biobank), showcasing its effectiveness in capturing brain dynamics across populations [19].
@@ 3.1 Experimental Settings
All experiments were carried out on the Perlmutter supercomputer, an HPE Cray EX system. A holdout strategy with a 70:15:15 data split was used for training, validation, and testing. To address data imbalance (see Section 2), stratified splits were applied based on the BSID-III scores. For benchmarking, 5-fold cross-validation ensured consistent and fair splits between experiments. For binary classification, we used weighted focal loss [26] to prioritize harder-to-classify (at-risk) cases. This approach significantly improved performance, especially for these challenging instances. Following the original SwiFT paper [19], we padded input images to 64x64x64 to align with the model architecture, adapting from the 96x96x96 padding used in the original work due to our native image size of 45x55x45. For classification tasks, we evaluated performance using the area under the ROC curve (AUC), accuracy, and balanced accuracy (ACCbal) to account for class imbalance. For regression tasks, mean absolute error (MAE) and mean squared error (MSE) were used, along with adjusted metrics (MAEadj and MSEadj) to reflect re-scaled values for interpretability.
@@ 3.2 Baselines
We evaluated the SwiFT model’s performance on fMRI data in an end-to-end manner, focusing on single-label and multi-label prediction tasks. Multi-label prediction was hypothesized to enhance performance due to strong correlations among Bayley scores (see Section 2), as supported by prior work [27]. The models were fine-tuned across a predefined hyperparameter search space, including data augmentation strategies (affine-only, intensity-only, and combined transformations), learning rates, batch sizes, and patch sizes. Sequence length of the input was varied to evaluate its impact: a baseline length of 20 was used per the original SwiFT paper [19], with additional experiments using lengths of 50 and 100 to assess temporal dynamics. Longer sequences have shown promise in prior studies for predicting intelligence in young adults and elders [19]. To leverage publicly available fMRI datasets, we pre-trained SwiFT using contrastive self-supervised learning [24] on data from the UK Biobank (UKB)[23] [22], Human Connectome Project (HCP)[20], and Adolescent Brain Cognitive Development (ABCD) study[21]. While this approach utilizes larger datasets, differences between neonatal and adult or adolescent brain data present potential limitations. To address size discrepancies, neonatal images were padded to 96x96x96, contrasting with experiments trained from scratch, where images were padded to 64x64x64 (see Section 3.1).
@@ 3.4 Dimensionality Reduction via Group ICA
Group Independent Component Analysis (Group ICA) [14] [15] is a widely used data-driven technique in neuroimaging that decomposes fMRI data into independent spatial and temporal components. These components represent distinct neural processes or functional networks. Group ICA reduces the dimensionality of high-dimensional fMRI data while preserving significant information and isolating biologically interpretable brain networks. In this study, we use Group ICA features as model inputs for both single-label and multi-label learning. These features encapsulate complex brain activity patterns in a compact representation, enhancing expressiveness and facilitating brain network-level analysis. We hypothesize that this approach will improve model interpretability and predictive performance by leveraging the reduced dimensionality and biologically relevant features.
@@ 3.4.1 Preparation
The workflow for Group ICA and functional connectivity analysis (see Figure 3) follows methodologies established for the UKB dataset [13] [12]. For the selected 100 subjects, we used MIGP (MELODIC’s Incremental Group-PCA) to extract the top 1200 weighted spatial eigenvectors, approximating traditional PCA while remaining computationally efficient for large datasets [16]. Spatial Independent Component Analysis (spatial-ICA) was then applied using FSL MELODIC [14] [15] to parcellate the data into spatially independent components (ICs). We evaluated dimensionalities of 25 and 100 ICs, consistent with prior work on the UKB dataset [13] [12] [17] [18]. Higher dimensionalities result in smaller, more numerous regions within spatial maps. Since FIX denoising was performed during preprocessing, we did not identify artifactual components, assuming their removal. Using the generated group ICA maps, we performed dual regression to obtain subject-specific spatial maps for each IC. Dual regression comprises two steps: (i) spatial regression, projecting group-level IC maps onto individual data, and (ii) temporal regression, deriving subject-specific timeseries for each IC. To further explore functional connectivity, we conducted seed-to-voxel (seed2voxel) analysis, correlating timeseries of specific brain regions with those of all other voxels. The resulting whole-brain functional connectivity maps served as input features for the SwiFT model to predict neurodevelopmental outcomes.
Figure 3: UML Flowchart depicting the workflow to extract features and connectivity maps by leveraging Group Independent Component Analysis, heavily inspired by [13][12] and [25].
@@ 3.4.2 Results
Discussion
This study demonstrates that integrating Group ICA-based dimensionality reduction with SwiFT significantly enhances predictions of neurodevelopmental outcomes from neonatal fMRI data. By extracting biologically meaningful features via ICA and leveraging multi-label learning, the approach preserves critical neural information while reducing computational complexity. SwiFT effectively models local and global spatiotemporal patterns in 4D fMRI data, underscoring the synergy between neuroscience-driven feature engineering and advanced machine learning. As can be seen in Table 3, our results indicate that ICA dimensionality (25 vs. 100 ICs) does not substantially impact performance, suggesting that ICA robustly captures essential neural features regardless of the number of components. Furthermore, single-label learning with ICA-extracted features outperformed raw fMRI predictions, particularly for language classification and motor regression tasks, highlighting ICA’s ability to retain biologically meaningful information while reducing noise. The combination of ICA and multi-label learning further enhanced predictive power, aligning with previous findings on the interrelated nature of developmental outcomes. These findings reinforce ICA’s value as a preprocessing step, enabling the model to focus on key neural networks and improve accuracy and efficiency. This targeted approach demonstrates the potential of integrating neuroscience-driven features with attention-based architectures for advancing neurodevelopmental research and clinical applications.
@@ 4 Interpretation via Integrated Gradients
As in the original SwiFT study [19], we used Integrated Gradient with Smoothgrad sQuare (IG-SQ), implemented through the Captum framework [28] [29] [30]. This method allowed us to identify spatial representations in brain regions that contribute to Bayley-III classification tasks. This analysis was based on the best-performing model, which combined multi-label prediction with ICA-extracted features. Leveraging ICs allowed for neuroscientifically meaningful interpretations by associating model outputs with known functional networks [40]. For each test sample, we generated 4D IG-SQ maps and excluded misclassified instances to ensure the validity of interpretations. The resultant maps were normalized, Gaussian-smoothed, and averaged across subjects to identify spatial patterns consistently associated with cognitive, language, and motor outcomes. Additionally, we aggregated these maps across ICs to derive comprehensive interpretations of brain activity relevant to each developmental domain.

### v3

@@ 1 Introduction
Brain development during the first few months of life is a period of rapid structural and functional reorganization, making it a critical window for identifying potential neurodevelopmental deficits. Accurately predicting developmental outcomes during this period is essential for enabling early interventions that can mitigate the lifelong impact of developmental delays. Neonatal fMRI data, such as those from the Developing Human Connectome Project (dHCP), have shown potential for predicting neurodevelopmental outcomes [1]. However, the spatiotemporal complexity of neonatal brain activity presents significant challenges for conventional analysis methods. This study investigates the potential of the Swin 4D fMRI Transformer (SwiFT) [15], a deep learning architecture designed to process high-dimensional fMRI data, for predicting neurodevelopmental outcomes from neonatal fMRI. Unlike existing methods, SwiFT leverages 4D spatiotemporal attention mechanisms to effectively capture dynamic brain connectivity patterns, offering a novel approach for analyzing neonatal fMRI data. Specifically, we aim to predict Bayley Scales of Infant and Toddler Development, Third Edition (Bayley-III/BSID-III) composite scores, encompassing cognitive, language, and motor skills, using neonatal fMRI from the dHCP dataset. To address the challenges of limited neonatal data and high dimensionality, we explore dimensionality reduction using group Independent Component Analysis (ICA) and pretraining SwiFT on large publicly available adult fMRI datasets. We hypothesize that (1) SwiFT outperforms baseline models in predicting Bayley-III scores due to its ability to capture complex spatiotemporal patterns in fMRI data and (2) incorporating ICA features and pretraining further improves predictive accuracy. Furthermore, we aim to interpret SwiFT’s predictions using Integrated Gradients with Smoothgrad sQuare (IG-SQ) [24] to identify brain regions associated with early cognitive, language, and motor development. This research advances the state of neurodevelopmental prediction by introducing a robust and interpretable framework for neonatal fMRI analysis. By addressing key methodological challenges, our findings lay the groundwork for future applications in neurodevelopmental research and clinical practice.
In Section 2, we present the dataset, its acquisition, processing, target values, and an overview of SwiFT. Section 3 outlines our experimental settings, including (i) baselines, (ii) SwiFT on raw fMRI data for single- and multi-label prediction, and (iii) SwiFT on group-ICA features for the same tasks. Section 4 focuses on IG-SQ analysis of the best model, interpreting IG maps for delays in cognitive, language, and motor skills. Finally, Section 5 discusses limitations and future directions.
@@ 2 Methods
Following prior studies [7], we simplify the classification task into a binary format, with a threshold of 85 distinguishing low- and high-risk cases. Regression tasks are also conducted to predict continuous Bayley-III scores.
@@
To predict developmental outcomes from neonatal fMRI, we used SwiFT, a Transformer-based model designed for high-dimensional spatiotemporal fMRI data. SwiFT processes 4D fMRI volumes by segmenting them into patches, embedding them as tokens, and passing them through layers of 4D Swin Transformer blocks. These blocks utilize a shifted window multi-head self-attention mechanism to capture both local and global spatiotemporal dependencies while minimizing computational costs through patch merging. SwiFT demonstrated robust performance on diverse tasks (e.g., predicting sex, age, intelligence, or psychiatric diagnosis), showcasing its effectiveness in learning spatiotemporal patterns from the human brains [15].
@@ 3.1 Experimental Settings
All experiments were carried out on the Perlmutter supercomputer, an HPE Cray EX system. A holdout strategy with a 70:15:15 data split was used for training, validation, and testing. To address data imbalance (see Section 2), stratified splits were applied based on the BSID-III scores. For benchmarking, 5-fold cross-validation ensured consistent and fair splits between experiments. For binary classification, we used weighted focal loss [22] to prioritize harder-to-classify (at-risk) cases. This approach significantly improved performance, especially for these challenging instances. Following  [15], we padded input images to 64x64x64 to align with the model architecture, adapting from the 96x96x96 padding used in  [15] due to our native image size of 45x55x45. For classification tasks, we evaluated performance using the area under the ROC curve (AUC), accuracy, and balanced accuracy (ACCbal) to account for class imbalance. For regression tasks, mean absolute error (MAE) and mean squared error (MSE) were used, along with adjusted metrics (MAEadj and MSEadj) to reflect re-scaled values for interpretability.
@@ 3.2 Baselines
We evaluated the SwiFT model’s performance on fMRI data in an end-to-end manner, focusing on single-label and multi-label prediction tasks. Multi-label prediction was hypothesized to enhance performance due to strong correlations among Bayley scores (see Section 2), as supported by prior work [23]. The models were fine-tuned across a predefined hyperparameter search space, including data augmentation strategies (affine-only, intensity-only, and combined transformations), learning rates, batch sizes, and patch sizes. Sequence length of the input was varied to evaluate its impact: a baseline length of 20 was used per  [15], with additional experiments using lengths of 50 and 100 to assess temporal dynamics, as longer sequences have shown promise in prior studies for predicting intelligence in young adults and elders [15]. To leverage publicly available fMRI datasets, we pre-trained SwiFT using contrastive self-supervised learning [20] on data from the UK Biobank (UKB) [19] [18], Human Connectome Project (HCP) [16], and Adolescent Brain Cognitive Development (ABCD) study [17]. While this approach utilizes larger datasets, differences between neonatal and adult or adolescent brain data present potential limitations. To address size discrepancies, neonatal images were padded to 96x96x96, contrasting with experiments trained from scratch, where images were padded to 64x64x64 (see Section 3.1).
@@ 3.4 Dimensionality Reduction via Group ICA
Group Independent Component Analysis (Group ICA) [10] [11] is a widely used data-driven technique in neuroimaging that decomposes fMRI data into independent spatial and temporal components. These components represent distinct neural processes or functional networks. Group ICA reduces the dimensionality of high-dimensional fMRI data while preserving significant information and isolating biologically interpretable brain networks. In this study, we use Group ICA features as inputs. These features encapsulate complex brain activity patterns in a compact representation, enhancing expressiveness and facilitating brain network-level analysis. We hypothesize that this approach will improve model interpretability and predictive performance by leveraging the reduced dimensionality and biologically relevant features.
@@ 3.4.1 Preparation
The workflow for Group ICA and functional connectivity analysis (see Figure 3) follows methodologies established for the UKB dataset [9] [8]. For the selected 100 subjects, we used MIGP (MELODIC’s Incremental Group-PCA) to extract the top 1200 weighted spatial eigenvectors, approximating traditional PCA while remaining computationally efficient for large datasets [12]. Spatial Independent Component Analysis (spatial-ICA) was then applied using FSL MELODIC [10] [11] to parcellate the data into spatially independent components (ICs). We evaluated dimensionalities of 25 and 100 ICs, consistent with prior work on the UKB dataset [9] [8] [13] [14]. Higher dimensionalities result in smaller, more numerous regions within spatial maps. Since FIX denoising was performed during preprocessing, we did not identify artifactual components. Using the generated group ICA maps, we performed dual regression to obtain subject-specific spatial maps for each IC. Dual regression comprises two steps: (i) spatial regression, projecting group-level IC maps onto individual data, and (ii) temporal regression, deriving subject-specific timeseries for each IC. To further explore functional connectivity, we conducted seed-to-voxel analysis, correlating timeseries of brain regions with those of all other voxels. The resulting whole-brain functional connectivity maps served as input features for the SwiFT model to predict neurodevelopmental outcomes.
Figure 3: UML Flowchart depicting the workflow to extract features and connectivity maps by leveraging Group Independent Component Analysis  [9][8] and [21].
@@ 3.4.2 Results
Discussion
This study demonstrates that integrating Group ICA-based dimensionality reduction with SwiFT significantly enhances predictions of neurodevelopmental outcomes from neonatal fMRI data. By extracting biologically meaningful features via ICA and leveraging multi-label learning, the approach preserves critical neural information while reducing computational complexity. SwiFT effectively learns local and global spatiotemporal patterns in 4D fMRI data, underscoring the synergy between neuroscience-driven feature engineering and advanced machine learning. As can be seen in Table 3, our results indicate that ICA dimensionality (25 vs. 100 ICs) does not substantially impact performance, suggesting that ICA robustly captures essential neural features regardless of the number of components. Furthermore, single-label learning with ICA-extracted features outperformed raw fMRI predictions, particularly for language classification and motor regression tasks, highlighting ICA’s ability to retain biologically meaningful information while reducing noise. The combination of ICA and multi-label learning further enhanced predictive power, aligning with previous findings on the interrelated nature of developmental outcomes. These findings reinforce ICA’s value as a preprocessing step, enabling the model to focus on key neural networks and improve accuracy and efficiency. This targeted approach demonstrates the potential of integrating neuroscience-driven features with attention-based architectures for advancing neurodevelopmental research.
@@ 4 Interpretation via Integrated Gradients
As in  [15], we used Integrated Gradient with Smoothgrad sQuare (IG-SQ), implemented through the Captum framework [24]. This method allowed us to identify spatial representations in brain regions that contribute to Bayley-III classification tasks. This analysis was based on the best-performing model, which combined multi-label prediction with ICA-extracted features. Leveraging ICs allowed for neuroscientifically meaningful interpretations by associating model outputs with known functional networks [34]. For each test sample, we generated 4D IG-SQ maps and excluded misclassified instances to ensure the validity of interpretations. The resultant maps were normalized, Gaussian-smoothed, and averaged across subjects to identify spatial patterns consistently associated with cognitive, language, and motor outcomes. Additionally, we aggregated these maps across ICs to derive comprehensive interpretations of brain activity relevant to each developmental domain.
@@ 5 Concluding Remarks
Acknowledgements
This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. 2021R1C1C1006503, RS-2023-00266787, RS-2023-00265406, RS-202400421268), by Creative-Pioneering Researchers Program through Seoul National University(No. 20020240057), by Semi-Supervised Learning Research Grant by SAMSUNG(No.A0426-20220118), by Identify the network of brain preparation steps for concentration Research Grant by LooxidLabs(No.33920230001), by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)] by the MSIT(Ministry of Science, ICT), Korea, under the Global Research Support Program in the Digital Field program(RS-2024-00421268) supervised by the IITP(Institute for Information & Communications Technology Planning & Evaluation), by the National Supercomputing Center with supercomputing resources including technical support(KSC-2023CRE-0568) and by the Ministry of Education of the Republic of Korea and the National Research Foundation of Korea (NRF - 2021S1A3A2A02090597), and by Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT(MSIT, Korea) & Gwangju Metropolitan City. This research used resources of: the Oak Ridge Leadership Computing Facility, a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725; the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility using NERSC award NERSC DDR-ERCAP0030592 and DDR-GenAI, and ALCC-ERCAP0030659.